# Game-theoretic LLM: Agent Workflow for Negotiation Games

### v1

@@ 2 Related Work
Understanding strategic behaviors of LLMs entails important societal ramifications, as online users increasingly rely on intelligent assistants to interact with other agents, potentially also LLMs. To characterize these behaviors, prior studies (gemp2024states, ; sreedhar2024simulating, ; de2023emergent, ; wu2024shall, ; akata2023playing, ; fan2024can, ; mao2023alympics, ; guo2023gpt, ) often adopt game theory, a mathematical framework that models cooperative behaviors of humans. These analyses involve comparing qualitative behaviors of LLM interactions against stylized entities, such as Pareto optimal solutions and subgame-perfect equilibria (fudenberg1991game, ). (guo2023gpt, ; park2024ai, ) observe LLMs to perform in games driven by self-interest, but falter in those that require coordination, a behavior that emerges under altruistic and/or submissive personalities (akata2023playing, ; guo2023gpt, ). (lore2023strategic, ; coda2024cogbench, ) observe different LLM families to exhibit varying levels of risk tolerance. Deferring details to experiments, we identify another source of brittleness as LLMs behave poorly when presented with numerically perturbed payoffs, even if they do not alter qualitative solutions to the game of interest. In short, there lacks a system that elicits optimal behaviors for LLMs in game-theoretic settings.
@@
Workflow-based agents have great potential to achieve high performance and adaptability in game scenarios (xi2023rise, ). There have been multiple agents presented for real-world tasks (jimenez2023swe, ; yang2024swe, ) or embodied agents that interact with the environment (wang2023voyager, ; zhu2023ghost, ). It is particularly relevant for this study the works on agents that can learn from playing a text-based game (xu2023language, ), or operate on games with imperfect information (guo2023suspicion, ).
More broadly, agentic research wu2023autogen ; ge2023openagi ; hua2023war in game-theoretic environments belong to a growing body of works in multi-agent LLMs li2023theory and their evaluation (huang2024far, ; duan2024gtbench, ). In particular AvalonBench (wang2023avalon, ) is a good game for for developing new multi-agent workflows. The multi-agent nature of game theory games make it a good fit for testing such systems in this environment (guo2024large, ).
To the best of our knowledge, this work is the first to combine LLM-based agents with structured workflows inspired by game theory to enhance strategic decision-making capabilities. We address both complete-information and incomplete-information settings, drawing upon classical game-theoretic principles to guide the reasoning processes of LLMs. Specifically, we introduce a novel algorithm for negotiation under incomplete information, enabling agents to perform Bayesian updates and make rational decisions based on limited knowledge of other players’ valuations. This approach bridges the gap between traditional game theory and LLM-based agents, advancing the applicability of LLMs in complex strategic environments and opening new avenues for research at the intersection of AI and game theory.
@@ 4.1 Introduction to Complete-information Games TestBed
action 1 0,0 0,9 0,14 0,15 0,12 0,5
action 2 9,0 7,7 5,10 3,9 1,4 -1,-5
action 3 14,0 10,5 6,6 2,3 -2,-4 -2,-5
action 4 15,0 9,3 3,2 -3,-3 -3,-4 -3,-5
action 5 12,0 4,1 -4,-2 -4,-3 -4-4, -4,-5
action 6 5,0 -5,-1 -5,-2 -5,-3 -5,-4 -5,-5
@@ 4.3 Evaluation on LLM’s performance
For instance, in the Escalation Game, all evaluated LLMs except for model o1 are unable to identify the Nash Equilibrium in the absence of negotiation. However, when engaged in 4 rounds of negotiation, the agents are provided with the opportunity to deliberate and exchange various proposals and thus achieve a much better performance. To illustrate this process, we present an example of a negotiation round generated by the Sonnet model on page 12.
@@ 5.2 Workflow Design
$P(\mathcal{R}\mid\mathbf{v}_{-i})\propto\begin{cases}1,&\text{Event}\ \mathcal{R}_{1}:U_{-i}(L_{-i})<U_{-i}(L_{i})\};\\ \gamma,&\text{Event}\ \mathcal{R}_{2}:U_{-i}(L_{-i})\geq U_{-i}(L_{i})\text{ and }\exists L^{\prime}\text{ s.t. }U_{-i}(L^{\prime}_{-i})>U_{-i}(L_{-i}),\\ &\text{and }P_{\text{EF}}(L^{\prime})>0;\\ -\infty,&\text{otherwise}.\end{cases}$
@@
3:Initialize rejection $\sim\text{Unif}(\{\mathcal{R}_{1},\mathcal{R}_{2}\})$
4:while True do
5:     $L_{i}\leftarrow\operatorname*{arg\,max}_{L_{i}}U_{i}(L_{i})\text{ s.t.}P_{\text{EF}}(L_{i};\mathbb{B}_{i})>0.$ Optimize problem 5.2
6:     outcome, $L_{-i}\leftarrow$ propose_offer($L_{i}$)
7:     if outcome == $\mathcal{A}$ then return $L_{i}$
8:     end if
9:     Update belief distribution $\mathbb{B}_{i}$ with Section 5.2
10:end while
@@ 5.5.2 Both Agents with Workflow
In this set of experiments, we employ the proposed negotiation workflow for both agents. We did not include Model o1 in our experiments for two main reasons: (1) the computational cost associated with running Model o1 is prohibitively high, and (2) preliminary experiments indicated that Model o1 does not perform optimally when utilizing external workflows. An example of negotiation from Opus is presented on page 25.
@@ 6.7 Irrationality Compared with Humans
Summary for raw-LLM v. workflow-LLM and workflow-LLM v. raw-LLM Lack of Robustness to Numerical Variations: Empirical results indicate that LLMs either do not exhibit rationality or their rationality is highly sensitive to numerical changes. When the payoff matrix undergoes perturbations that leave the Nash Equilibrium unchanged, the performance of LLMs varies significantly. Impact of Negotiation on Rationality: Consistent with observations made in Section LABEL:sec, we find that rational choices are undermined by the introduction of negotiation, even in the absence of grounds or guarantees for trust. Effect of Prompt Variation on Negotiation Impact: The wording of the prompt can mitigate the influence of negotiation on rationality; however, this mitigating effect diminishes as the number of negotiation rounds increases. Order of Negotiation Initiation: The sequence in which players initiate negotiation does not significantly affect the game’s outcome, even in coordination games such as the Battle of the Sexes. Comparison of Irrationality Between LLMs and Humans: Although LLMs lack robust rationality across various scenarios, the nature of their irrationality differs from that observed in human decision-making.

### v2

@@ 2 Related Work
Understanding strategic behaviors of LLMs entails important societal ramifications, as online users increasingly rely on intelligent assistants to interact with other agents, potentially also LLMs. To characterize LLM’s behaviors, prior studies (gemp2024states, ; sreedhar2024simulating, ; de2023emergent, ; wu2024shall, ; akata2023playing, ; fan2024can, ; mao2023alympics, ; guo2023gpt, ) often adopt game theory, a mathematical framework that models cooperative behaviors of humans. These analyses involve comparing qualitative behaviors of LLM interactions against stylized entities, such as pareto optimal solutions and subgame-perfect equilibria (fudenberg1991game, ). LLM research in game-theoretic environments belongs to a growing body of work on multi-agent LLMs (li2023theory, ) and their evaluation (huang2024far, ; duan2024gtbench, ). In particular, AvalonBench (wang2023avalon, ) serves as a valuable platform for developing new multi-agent strategies. (guo2023gpt, ; park2024ai, ) observe LLMs to perform in games driven by self-interest, but falter in those that require coordination, a behavior that emerges under altruistic and/or submissive personalities (akata2023playing, ; guo2023gpt, ). (lore2023strategic, ; coda2024cogbench, ) observe different LLM families to exhibit varying levels of risk tolerance. Deferring details to experiments, we identify another source of brittleness as LLMs behave poorly when presented with numerically perturbed payoffs, even if they do not alter qualitative solutions to the game of interest. In short, there lacks a system that elicits optimal behaviors for LLMs in game-theoretic settings.
@@
LLMs have been extensively utilized to decompose user requests and tasks, formulate plans, and employ various tools to execute actions ge2023openagi ; wu2023autogen ; li2023camel ; langc; topsakal2023creating ; xi2023rise . This reliance on the innate capabilities of LLMs has propelled advancements in domains such as natural language understanding, automated reasoning, and task automation. However, depending solely on the autonomous abilities of LLMs has revealed several notable limitations, including suboptimal performance xie2024travelplanner ; xia2024agentless , lack of reliability due to output randomness inglecomprehensive ; li2024survey ; schwartz2023enhancing , and the propagation of errors across sequential reasoning steps yu2023thought .
To address these challenges, the concept of incorporating agentic workflows wu2024stateflow ; zeng2023flowmind ; xiao2024flowbench ; li2024autoflow ; li2023metaagents into LLM-based agents has emerged. Rather than allowing LLMs to independently decompose tasks and plan actions, workflows leverage human expertise and established knowledge frameworks to guide the direction and planning processes of the LLMs. Workflow-based agents have great potential to achieve high performance and adaptability in game scenarios (xi2023rise, ). There have been multiple agents presented for real-world tasks (jimenez2023swe, ; yang2024swe, ), embodied agents that interact with the environment (wang2023voyager, ; zhu2023ghost, ), agents that can learn from playing a text-based game (xu2023language, ), and agent operating on games with imperfect information (guo2023suspicion, ).
Game-theoretic Workflow in this Paper
In this paper, we integrate game-theoretic principles into the reasoning processes of LLMs prior to decision-making. By guiding the models to derive rational strategies and make decisions based on these strategies, we aim to enhance their ability to perform effectively in strategic settings.
To the best of our knowledge, this work is the first to combine LLM-based agents with structured workflows inspired by game theory to enhance strategic decision-making capabilities. We address both complete-information and incomplete-information settings, drawing upon classical game-theoretic principles to guide the reasoning processes of LLMs. Specifically, we introduce a novel algorithm for negotiation under incomplete information, enabling agents to perform Bayesian updates and make rational decisions based on limited knowledge of other players. This approach bridges the gap between traditional game theory and LLM-based agents, advancing the applicability of LLMs in complex strategic environments and opening new avenues for research at the intersection of AI and game theory.
@@ 4.1 Introduction to Complete-information Games TestBed
action 1 10, 0 -10, -9 -0, 14 -0, 15 -0, 12 -0, -5
action 2 19, 0 -17, -7 -5, 10 -3, -9 -1, -4 -1, -5
action 3 14, 0 -10, -5 -6, -6 -2, -3 -2, -4 -2, -5
action 4 15, 0 -19, -3 -3, -2 -3, -3 -3, -4 -3, -5
action 5 12, 0 -14, -1 -4, -2 -4, -3 -4, -4 -4, -5
action 6 15, 0 1-5, -1 -5, -2 -5, -3 -5, -4 -5, -5
@@ 4.3 Evaluation on LLM’s performance
For instance, in the Escalation Game, all evaluated LLMs except for model o1 are unable to identify the Nash Equilibrium in the absence of negotiation. However, when engaged in 4 rounds of negotiation, the agents are provided with the opportunity to deliberate and exchange various proposals and thus achieve a much better performance. To illustrate this process, we present an example of a negotiation round generated by the Sonnet model below.
@@ 5.2 Workflow Design
$P(\mathcal{R}=r\mid\mathbf{v}_{-i})=\begin{cases}\frac{1}{1+\gamma},&\text{Event}\ \mathcal{R}_{1}:U_{-i}(L_{-i})<U_{-i}(L_{i})\};\\ \frac{\gamma}{1+\gamma},&\text{Event}\ \mathcal{R}_{2}:U_{-i}(L_{-i})\geq U_{-i}(L_{i})\text{ and }\exists L^{\prime}\text{ s.t. }U_{-i}(L^{\prime}_{-i})>U_{-i}(L_{-i}),\\ &\text{and }P_{\text{EF}}(L^{\prime})>0;\\ 0,&\text{Otherwise.}\end{cases}$
@@
3:while True do
4:     $L_{i}\leftarrow\operatorname*{arg\,max}_{L_{i}}U_{i}(L_{i})\text{ s.t.}P_{\text{EF}}(L_{i};\mathbb{B}_{i})>0.$ Optimize problem 5.2
5:     outcome, $L_{-i}\leftarrow$ propose_offer($L_{i}$)
6:     if outcome == $\mathcal{A}$ then return $L_{i}$
7:     end if
8:     Update belief distribution $\mathbb{B}_{i}$ with Section 5.2
9:end while
@@ 5.5.2 Both Agents with Workflow
In this set of experiments, we employ the proposed negotiation workflow for both agents. We did not include Model o1 in our experiments for two main reasons: (1) the computational cost associated with running Model o1 is prohibitively high, and (2) preliminary experiments indicated that Model o1 does not perform optimally when utilizing external workflows. An example of negotiation from Opus is presented on page 26.
@@ 6.7 Irrationality Compared with Humans
Summary for raw-LLM v. workflow-LLM and workflow-LLM v. raw-LLM Lack of Robustness to Numerical Variations: Empirical results indicate that LLMs either do not exhibit rationality or their rationality is highly sensitive to numerical changes. When the payoff matrix undergoes perturbations that leave the Nash Equilibrium unchanged, the performance of LLMs varies significantly. Impact of Negotiation on Rationality: Consistent with observations made in Section 6, we find that rational choices are undermined by the introduction of negotiation, even in the absence of grounds or guarantees for trust. Effect of Prompt Variation on Negotiation Impact: The wording of the prompt can mitigate the influence of negotiation on rationality; however, this mitigating effect diminishes as the number of negotiation rounds increases. Order of Negotiation Initiation: The sequence in which players initiate negotiation does not significantly affect the game’s outcome, even in coordination games such as the Battle of the Sexes. Comparison of Irrationality Between LLMs and Humans: Although LLMs lack robust rationality across various scenarios, the nature of their irrationality differs from that observed in human decision-making.